OpenStudy (anonymous):

solve the equation for the unknown (5x)/2-7/8=14

OpenStudy (mimi_x3):

\[\frac{5x}{2} -\frac{7}{8} = 14\] \[\frac{5x\times4}{2\times4} -\frac{7}{8} =14\] \[\frac{20x}{8} -\frac{7}{8}=14 \] \[\frac{20x-7}{8} =14\] \[20x-7=112 \] \[20x = 112+7 \] \[x=...\]

OpenStudy (mimi_x3):

You work that out yourself. At least try.

OpenStudy (mimi_x3):

Yes!